OpenStudy (anonymous):

A BULLET OF MASS a AND b IS FIRED INTO A LARGE BLOCK OF MASS c AND STICK TO IT. THE VELOCITY OF THE SYSTEM AFTER THE COLLISION IS

OpenStudy (anonymous):

Here we can use conservation of momentum (p=mv): (a+b)vi = (a+b+c)vf vf = (a+b)vi/(a+b+c)